OpenStudy (anonymous):

60 pens in 25 boxes 12 pens in y boxes

OpenStudy (anonymous):

This is proportional so: \[\frac{ 60 pens }{ 25boxes } = \frac{ 12pens }{ yboxes }\] Then you would just cross multiply: 60y = 300 Divide both sides by 60: y = 300 / 60 y = 5 boxes

OpenStudy (arabpride):

yerp ~ setting up the proportion iz exactly how u would've solved this one *^.^*

OpenStudy (anonymous):

SpooksDF is correct